1. Methylhydrazine (N2H3CH3) is a common liquid propellant used in rocket fuels. Look for the standard molar enthalpies of formation of the reactants and products (you may use the uploaded files or search elsewhere for this). Calculate the ∆Ho for the reaction.
  • N2H3CH3 (l) + N2O4 (l) ---> H2(g) + N2 (g) + CO2 (g)
1
Expert's answer
2021-10-20T02:51:19-0400

Solution.

"Q(cal.) = Q(reac.)"

"Q(cal.) = C(cal.) × \Delta T = 113.013 kJ"

"Q(reac/mol) = \frac{Q(reac.)}{n(CH6N2)}"

"n(CH6N2) = \frac{m(CH6N2)}{M(CH6N2)}"

"n(CH6N2) = \frac{4.00}{46.1} = 0.09 \ mol"

"Q(reac/mol) = \frac{113.013 \times 10^3}{0.09} = 1.30 \times 10^6 \ J/mol"

Answer:

"Q(reac/mol) = 1.30 \times 10^6 \ J/mol"